OpenStudy (alyssajobug):

Solve the linear equasion y=-2x-8 and y=6x+1

OpenStudy (anonymous):

what exactly are u looking for

OpenStudy (marigirl):

I guess you are trying to find the value of x in which both the equations will be the same or intercept? you can make the equations equal to each other \[-2x-8=6x+1\] \[-2x-6x=1+8\] \[-8x=9\] \[x=\frac{ - 9}8{ ? }\]

OpenStudy (alyssajobug):

Write the solution as an ordered pair like this: (10, -7)

OpenStudy (wolf1728):

First put the equations in standard form: A) 2x + y = -8 B) -6x +y = +1 Multiply equation B by minus 1 B) 6x -y = -1 then add this to A) A) 2x + y = -8 8x = -9 x = -1.125 Inputting this value into equation A 2*-1.125 +y = -8 -2.25 +y = -8 y = -5.75

OpenStudy (mathstudent55):

I think you mean "solve the system of linear equations." y = -2x - 8 y = 6x + 1 Subtract the second equation from the first equation: 0 = -8x - 9 Add 8x to both sides 8x = -9 Divide both sides by 8 \(x = -\dfrac{9}{8} \) Now that we know the value of x, substitute x in the first equation with the klnown value. y = -2x - 8 \(y = -2 \cdot \left( -\dfrac{9}{8} \right) - 8 \) Write 2 as a fraction to multiply \(y = -\dfrac{2}{1} \cdot \left( -\dfrac{9}{8} \right) - 8 \) Multiply the fractions \(y = \dfrac{18}{8} - 8 \) Reduce the fraction \(y = \dfrac{9}{4} - 8 \) Write 8 as a fraction with denominator 4 to be able to add \(y = \dfrac{9}{4} - \dfrac{32}{4} \) Add fractions \(y = -\dfrac{23}{4}\) The solution is \( \left( -\dfrac{9}{8}, -\dfrac{23}{4} \right) \)